Assume that the length of a phone call (minutes) is a continuous random variable X with probability density function {f(x)=1/10e^-0.1x} . If someone arrives at a phone booth just before you arrive, find the probability that you have to visit


1). Less the 5 minutes


2). Greater than 5 minutes


3). Between 5 and 20 minutes

1
Expert's answer
2021-12-28T16:26:15-0500

We are given that "f(x) = 0.1e ^{\u22120.1x} , x \u2265 0" .

1)


"P(X<5)=\displaystyle\int_{0}^50.1e ^{−0.1x}dx=[-e ^{−0.1x}]\begin{matrix}
 5 \\
 0
\end{matrix}"

"=1-e^{-0.5}\approx0.3935"

2)


"P(X>5)=\displaystyle\int_{5}^{\infin}0.1e ^{−0.1x}dx=[-e ^{−0.1x}]\begin{matrix}
 \infin \\
 5
\end{matrix}"

"=e^{-0.5}\approx0.6065"

3)


"P(5<X<20)=\displaystyle\int_{5}^{20}0.1e ^{−0.1x}dx=[-e ^{−0.1x}]\begin{matrix}
 20 \\
 5
\end{matrix}"

"=e^{-0.5}-e^{-2}\approx0.4712"